# Generative Autoregressive Transformers for Model-Agnostic Federated MRI Reconstruction

### v1

@@ 4.3 Modeling Procedures
We considered several three-site FL setups to examine performance under model-heterogeneous settings. The reconstruction models trained in these setups included MoDL [14], rGAN [20], and D5C5 [13]. All sites were taken to use common imaging operators with matched acceleration rates and k-space sampling densities. Each site was allowed to train a distinct reconstruction model. In a first experiment, the three sites trained entirely distinct architectures (MoDL-3, rGAN, D5C5). In a second experiment, the sites trained the same model type albeit using different numbers of cascades (MoDL-3, MoDL-5, MoDL-7). All models were implemented using the PyTorch framework and executed on RTX 4090 GPUs. Models were trained via the AdamW optimizer with $\beta_{1}$=\,=\,=0.9, $\beta_{2}$=\,=\,=0.95, a weight decay of 0.05, and a learning rate of $\eta$=\,=\,=$10^{-4}$.
As model-heterogeneous settings preclude aggregation via simple averaging of model weights, FL baselines used either model distillation or synthetic image generation as a means for cross-site knowledge transfer. Note that distillation-based aggregation involves bidirectional distillation between global and site-specific models, with each direction incurring significantly higher computational costs compared to model averaging, as the student model must be trained based on a sizable set of input-output images produced by the teacher model. Here, we observed that distillation-based baselines (FedDF, FedMD) attained a near-optimal trade-off between performance and efficiency with a total of 200 training epochs and distillation every 100 epochs. In generative baselines (FedGIMP, FedDDA), the two-tier strategy in FedGAT was followed under matching training procedures with $N_{c}$=\,=\,=500, $N_{l}$=\,=\,=1, $N_{p}$=\,=\,=100 and $N_{f}$=\,=\,=100.
@@
Figure 3: Representative reconstructions at R=8x from zero-filled Fourier method (Zero-filled), single-site models (Single), FL baselines (FedDF, FedMD, FedGIMP, FedDDA), and FedGAT, along with reference images. (a) Site-specific BraTS model tested on fastMRI, (b) Site-specific fastMRI model tested on MIDAS, (c) Site-specific MIDAS model on BraTS. Zoom-in displays are included showing error maps (left) and images (right) within a focal window to emphasize method differences.
Table 3: Within-site and cross-site reconstruction performance for single-coil datasets, where each site employs a variant of MoDL with different number of cascades.
Within-site Site 1 (BraTS-MoDL-5) Site 2 (fastMRI-MoDL-3) Site 3 (MIDAS-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 51.80$\pm$2.83 99.86$\pm$0.06 40.79$\pm$2.94 97.09$\pm$1.26 36.31$\pm$2.89 95.91$\pm$1.71
8x 44.07$\pm$3.26 99.38$\pm$0.33 36.30$\pm$2.83 95.14$\pm$1.81 30.93$\pm$2.81 92.45$\pm$3.01
Central 4x 51.30$\pm$2.38 99.83$\pm$0.08 40.00$\pm$3.12 96.76$\pm$1.36 36.06$\pm$2.76 95.73$\pm$1.73
8x 44.19$\pm$2.79 99.26$\pm$0.36 35.02$\pm$3.00 94.14$\pm$2.05 31.08$\pm$2.80 92.47$\pm$2.97
FL methods FedDF 4x 47.08$\pm$2.88 99.63$\pm$0.13 39.31$\pm$3.12 96.35$\pm$1.55 33.31$\pm$2.71 94.19$\pm$2.13
8x 41.53$\pm$3.13 98.89$\pm$0.46 34.18$\pm$3.13 93.49$\pm$2.48 29.54$\pm$2.55 90.28$\pm$3.28
FedMD 4x 47.02$\pm$2.84 99.63$\pm$0.13 39.23$\pm$3.27 96.37$\pm$1.58 33.37$\pm$2.69 94.20$\pm$2.12
8x 41.43$\pm$3.10 98.86$\pm$0.47 33.97$\pm$3.16 93.52$\pm$2.46 30.03$\pm$2.53 90.52$\pm$3.25
FedGIMP 4x 46.33$\pm$3.07 99.56$\pm$0.14 39.46$\pm$3.60 96.38$\pm$1.78 34.71$\pm$2.88 94.66$\pm$2.21
8x 41.60$\pm$2.89 98.72$\pm$0.42 34.90$\pm$3.32 93.97$\pm$2.57 30.28$\pm$2.78 90.56$\pm$3.63
FedDDA 4x 47.38$\pm$2.81 99.66$\pm$0.12 39.22$\pm$3.35 96.53$\pm$1.51 29.16$\pm$1.39 60.74$\pm$2.14
8x 42.92$\pm$2.99 99.08$\pm$0.40 34.14$\pm$3.36 93.86$\pm$2.43 27.80$\pm$2.03 63.50$\pm$2.74
FedGAT 4x 50.42$\pm$2.44 99.81$\pm$0.08 40.01$\pm$3.20 96.78$\pm$1.35 35.33$\pm$2.90 95.57$\pm$1.72
8x 43.43$\pm$2.85 99.25$\pm$0.35 35.50$\pm$2.91 94.54$\pm$2.10 30.04$\pm$2.87 91.92$\pm$3.10
Across-site Site 1 (BraTS-MoDL-5) Site 2 (fastMRI-MoDL-3) Site 3 (MIDAS-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 34.21$\pm$4.48 90.92$\pm$4.36 39.25$\pm$6.94 95.83$\pm$4.14 45.38$\pm$5.04 98.28$\pm$1.77
8x 30.79$\pm$4.08 87.83$\pm$5.52 34.46$\pm$6.48 92.76$\pm$5.57 38.99$\pm$4.18 97.09$\pm$2.44
Central 4x 38.98$\pm$4.55 96.36$\pm$1.80 40.83$\pm$7.05 96.78$\pm$3.19 47.23$\pm$5.35 98.62$\pm$1.50
8x 33.82$\pm$4.45 93.52$\pm$3.15 35.31$\pm$6.20 93.43$\pm$5.26 41.77$\pm$4.41 97.75$\pm$2.03
FL methods FedDF 4x 36.36$\pm$2.96 95.21$\pm$1.92 39.33$\pm$2.66 96.23$\pm$1.74 43.44$\pm$3.03 98.09$\pm$1.04
8x 31.94$\pm$2.87 92.09$\pm$2.86 34.91$\pm$2.80 93.62$\pm$2.48 37.86$\pm$3.15 96.31$\pm$1.75
FedMD 4x 36.28$\pm$2.92 95.24$\pm$1.87 39.55$\pm$2.75 96.28$\pm$1.79 43.43$\pm$3.02 98.08$\pm$1.04
8x 31.91$\pm$2.85 92.16$\pm$2.81 34.87$\pm$2.84 93.63$\pm$2.53 37.58$\pm$3.25 96.37$\pm$1.72
FedGIMP 4x 36.97$\pm$4.56 94.87$\pm$3.09 38.71$\pm$6.51 95.25$\pm$4.62 44.10$\pm$4.08 98.33$\pm$1.74
8x 32.89$\pm$4.50 92.33$\pm$3.89 34.65$\pm$6.20 92.31$\pm$5.76 39.02$\pm$3.74 96.84$\pm$2.79
FedDDA 4x 37.52$\pm$4.02 95.48$\pm$2.51 39.81$\pm$6.80 95.91$\pm$4.06 32.27$\pm$1.40 59.96$\pm$6.01
8x 32.98$\pm$4.21 92.71$\pm$3.53 35.06$\pm$6.15 93.16$\pm$5.36 31.45$\pm$1.78 59.85$\pm$7.96
FedGAT 4x 38.07$\pm$4.48 96.09$\pm$1.93 40.32$\pm$7.04 96.33$\pm$3.66 46.68$\pm$5.21 98.68$\pm$1.65
8x 33.35$\pm$4.23 93.16$\pm$3.37 35.13$\pm$6.94 93.64$\pm$5.43 40.65$\pm$4.37 97.42$\pm$2.31
Table 4: Within-site and cross-site reconstruction performance for multi-coil datasets, where each site employs a variant of MoDL with different number of cascades.
Within-site Site 1 (f.knee-MoDL-3) Site 2 (f.brain-MoDL-5) Site 3 (UMRAM-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 33.78$\pm$2.27 93.37$\pm$1.52 40.33$\pm$3.45 98.03$\pm$0.84 49.96$\pm$2.06 99.61$\pm$0.17
8x 29.39$\pm$2.06 83.90$\pm$3.08 34.21$\pm$3.36 94.33$\pm$2.01 42.11$\pm$1.90 98.25$\pm$0.73
Central 4x 33.69$\pm$2.32 93.20$\pm$1.63 40.95$\pm$3.84 98.05$\pm$1.19 46.36$\pm$2.22 99.32$\pm$0.30
8x 29.39$\pm$2.19 83.62$\pm$3.09 34.27$\pm$3.50 94.11$\pm$2.28 39.73$\pm$1.86 97.87$\pm$0.72
FL methods FedDF 4x 32.33$\pm$2.36 90.47$\pm$2.44 38.87$\pm$3.62 95.66$\pm$1.49 44.56$\pm$2.72 98.52$\pm$0.63
8x 28.36$\pm$2.31 80.66$\pm$4.14 32.97$\pm$3.73 91.37$\pm$3.37 37.00$\pm$2.30 96.73$\pm$1.40
FedMD 4x 32.93$\pm$2.35 91.96$\pm$1.77 39.94$\pm$3.94 97.49$\pm$1.09 44.63$\pm$3.05 99.25$\pm$0.28
8x 28.15$\pm$2.42 81.60$\pm$3.52 33.85$\pm$3.79 93.68$\pm$2.55 37.55$\pm$2.55 97.09$\pm$1.20
FedGIMP 4x 33.78$\pm$2.27 93.41$\pm$1.48 40.54$\pm$3.38 97.88$\pm$0.90 44.99$\pm$2.09 99.02$\pm$0.32
8x 28.97$\pm$2.12 83.68$\pm$3.07 34.19$\pm$3.37 93.80$\pm$3.22 39.53$\pm$1.70 97.68$\pm$0.77
FedDDA 4x 32.53$\pm$2.33 91.75$\pm$1.82 39.60$\pm$3.62 97.52$\pm$1.38 42.57$\pm$2.92 97.32$\pm$0.78
8x 28.09$\pm$2.22 81.90$\pm$2.93 33.25$\pm$3.52 93.61$\pm$2.76 37.51$\pm$2.17 93.63$\pm$1.50
FedGAT 4x 34.10$\pm$2.24 93.44$\pm$1.46 40.87$\pm$3.58 98.03$\pm$1.27 45.13$\pm$2.14 99.22$\pm$0.29
8x 29.49$\pm$2.21 83.73$\pm$3.07 34.20$\pm$3.48 93.85$\pm$2.52 40.21$\pm$1.71 97.84$\pm$0.66
Across-site Site 1 (f.knee-MoDL-3) Site 2 (f.brain-MoDL-5) Site 3 (UMRAM-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 38.40$\pm$3.26 97.32$\pm$1.67 39.42$\pm$4.49 96.91$\pm$2.13 37.60$\pm$4.58 96.03$\pm$2.34
8x 32.39$\pm$3.04 92.63$\pm$2.94 32.99$\pm$3.57 90.55$\pm$6.10 31.48$\pm$3.96 88.07$\pm$5.35
Central 4x 39.70$\pm$3.36 97.62$\pm$1.70 40.64$\pm$4.84 97.36$\pm$1.99 40.39$\pm$4.11 97.53$\pm$1.31
8x 34.33$\pm$3.12 93.86$\pm$3.04 33.61$\pm$3.89 91.38$\pm$5.74 34.47$\pm$3.50 91.89$\pm$4.04
FL methods FedDF 4x 38.15$\pm$3.41 96.40$\pm$2.11 38.28$\pm$4.64 95.18$\pm$3.56 37.37$\pm$4.12 94.73$\pm$2.39
8x 32.94$\pm$3.12 92.43$\pm$3.47 32.69$\pm$4.11 89.39$\pm$7.34 30.15$\pm$3.64 86.84$\pm$5.61
FedMD 4x 38.75$\pm$3.44 97.01$\pm$1.71 39.32$\pm$4.78 96.80$\pm$2.28 38.39$\pm$4.49 95.92$\pm$2.28
8x 32.52$\pm$2.90 92.24$\pm$3.02 33.19$\pm$3.76 90.90$\pm$6.08 31.90$\pm$3.81 87.51$\pm$5.55
FedGIMP 4x 37.87$\pm$3.13 97.23$\pm$1.50 38.93$\pm$4.32 96.75$\pm$2.16 37.40$\pm$4.38 96.35$\pm$1.76
8x 32.01$\pm$2.98 92.52$\pm$2.91 32.96$\pm$3.63 90.71$\pm$5.94 31.70$\pm$4.13 89.29$\pm$5.18
FedDDA 4x 35.54$\pm$3.34 94.49$\pm$2.44 37.78$\pm$3.78 95.65$\pm$1.52 36.61$\pm$4.16 95.15$\pm$2.36
8x 30.15$\pm$3.57 88.74$\pm$3.98 32.10$\pm$3.43 89.21$\pm$5.03 30.80$\pm$3.74 86.32$\pm$5.53
FedGAT 4x 39.25$\pm$3.09 97.50$\pm$1.52 39.65$\pm$4.53 97.16$\pm$2.07 39.16$\pm$4.33 97.10$\pm$1.60
8x 33.52$\pm$2.98 92.90$\pm$3.11 33.58$\pm$3.93 90.93$\pm$5.81 32.98$\pm$3.89 90.29$\pm$4.57
@@ 4.4 Datasets
In single-coil experiments, multi-contrast coil-combined magnitude MR images from public BraTS [70], fastMRI [71] and MIDAS [72] datasets were analyzed. Each dataset was taken to correspond to a separate site. At each site, (training, validation, test) splits of (40, 5, 10) non-overlapping subjects were used, resulting in (840, 105, 210) cross-sectional images. Subjects were selected at random from the datasets.
@@
In multi-coil experiments, multi-contrast k-space data from public fastMRI-brain, fastMRI-knee [71] and an in-house dataset were analyzed [31]. Experimental procedures for collecting the in-house dataset were approved by the local ethics committee at Bilkent University, and written informed consent was obtained from all participants. Each dataset was taken to correspond to a separate site. At each site, (training, validation, test) splits of (36, 6, 8) non-overlapping subjects were used, resulting in (840, 105, 210) cross-sectional images. Subjects were selected at random from the datasets. Multi-coil data were compressed to 5 virtual coils [73].
@@
Both actual and synthetic k-space data were retrospectively undersampled using a variable-density (VD) pattern with acceleration factors R = 4x and 8x [1]. Models were trained on a mixture of multi-contrast MRI datasets without special procedures to handle separate contrasts. When necessary, actual and synthetic MR images were zero-padded to a consistent size to permit training of reconstruction models, albeit performance assessments were conducted in native image sizes by discarding the padded regions following reconstruction.

### v2

@@ 4.3 Modeling Procedures
We considered several three-site FL setups to examine performance under model-heterogeneous settings. The reconstruction models trained in these setups included MoDL [13], rGAN [19], and D5C5 [12]. All sites used a common acceleration rate and k-space sampling density. Each site was allowed to train a distinct reconstruction model. In a first experiment, the three sites trained entirely distinct architectures based on adversarial or unrolled backbones (MoDL, rGAN, D5C5). In a second experiment, the sites trained the same type of physics-driven model, albeit using different numbers of cascades in the unrolled backbone (MoDL3, MoDL5, MoDL7). In experiments involving a non-participating site, an unrolled TransUNet [79] architecture was used. All models were implemented using the PyTorch framework and executed on RTX 4090 GPUs. Models were trained via the AdamW optimizer with $\beta_{1}$=\,=\,=0.9, $\beta_{2}$=\,=\,=0.95, a weight decay of 0.05, and a learning rate of $\eta$=\,=\,=$10^{-4}$.
As model-heterogeneous settings preclude aggregation via simple weight averaging, FL baselines used either model distillation or synthetic image generation as a means for cross-site knowledge transfer. Note that distillation-based aggregation involves bidirectional distillation between global and site-specific models, with each direction incurring significantly higher computational costs compared to model averaging, as the student model must be trained based on a sizable set of input-output images produced by the teacher model. Here, we observed that distillation-based baselines (FedDF, FedMD) attained a near-optimal trade-off between performance and efficiency with a total of 200 training epochs and distillation every 100 epochs. In generative baselines (FedGIMP, FedDDA), the two-tier strategy in FedGAT was followed under matching training procedures with $N_{c}$=\,=\,=500, $N_{l}$=\,=\,=1, $N_{p}$=\,=\,=100 and $N_{f}$=\,=\,=100.
@@ 4.4 Datasets
In coil-combined experiments, multi-contrast coil-combined magnitude MR images from public BraTS [80], fastMRI [81] and MIDAS [82] datasets were analyzed. Each dataset was taken to correspond to a separate site. At each site, (training, validation, test) splits of (40, 5, 10) non-overlapping subjects were used, resulting in (840, 105, 210) cross-sectional images. Subjects were selected at random from the datasets.
@@
In multi-coil experiments, multi-contrast k-space data from public fastMRI-brain, fastMRI-knee [81] and an in-house dataset were analyzed [31]. Experimental procedures for collecting the in-house dataset were approved by the local ethics committee at Bilkent University, and written informed consent was obtained from all participants. In experiments involving a non-participating site, multi-coil k-space data from the Calgary dataset [77] were analyzed. Each dataset was taken to correspond to a separate site. At each site, (training, validation, test) splits of (36, 6, 8) non-overlapping subjects were used, resulting in (840, 105, 210) cross-sectional images. Subjects were selected at random from the datasets. Multi-coil data were compressed to 5 virtual coils [83].
@@
Both actual and synthetic k-space data were retrospectively undersampled using a variable-density (VD) pattern with acceleration factors R = 4x-8x [1]. Models were trained on a mixture of multi-contrast MRI datasets. When necessary, actual and synthetic MR images were zero-padded to a consistent size to permit training of reconstruction models, and performance assessments were conducted in native image sizes by discarding the padded regions following reconstruction.
Figure 3: Representative reconstructions at R=8x from zero-filled Fourier method (Zero-filled), single-site models (Single), FL baselines (FedDF, FedMD, FedGIMP, FedDDA), and FedGAT, along with reference images. (a) Site-specific BraTS model tested on fastMRI, (b) Site-specific fastMRI model tested on MIDAS, (c) Site-specific MIDAS model tested on BraTS. Zoom-in displays are included showing error maps (left; see colorbar) and images (right) within a focal window to emphasize method differences.
Table 3: Within-site and across-site reconstruction performance for coil-combined datasets (Site 1: BraTS, Site 2: fastMRI, Site 3: MIDAS). Each site employs an MoDL variant with different number of cascades.
Within-site Site 1 (MoDL5) Site 2 (MoDL3) Site 3 (MoDL7) Average
[…]
Non-FL Single 4x 51.8$\pm$2.8 99.9$\pm$0.1 40.8$\pm$2.9 97.1$\pm$1.3 36.3$\pm$2.9 95.9$\pm$1.7 43.0 97.6
8x 44.1$\pm$3.3 99.4$\pm$0.3 36.3$\pm$2.8 95.1$\pm$1.8 30.9$\pm$2.8 92.5$\pm$3.0 37.1 95.7
Central 4x 51.3$\pm$2.4 99.8$\pm$0.1 40.0$\pm$3.1 96.8$\pm$1.4 36.1$\pm$2.8 95.7$\pm$1.7 42.5 97.4
8x 44.2$\pm$2.8 99.3$\pm$0.4 35.0$\pm$3.0 94.1$\pm$2.0 31.1$\pm$2.8 92.5$\pm$3.0 36.8 95.3
FL methods FedDF 4x 47.1$\pm$2.9 99.6$\pm$0.1 39.3$\pm$3.1 96.4$\pm$1.6 33.3$\pm$2.7 94.2$\pm$2.1 39.9 96.7
8x 41.5$\pm$3.1 98.9$\pm$0.5 34.2$\pm$3.1 93.5$\pm$2.5 29.5$\pm$2.6 90.3$\pm$3.3 35.1 94.2
FedMD 4x 47.0$\pm$2.8 99.6$\pm$0.1 39.2$\pm$3.3 96.4$\pm$1.6 33.4$\pm$2.7 94.2$\pm$2.1 39.9 96.7
8x 41.4$\pm$3.1 98.9$\pm$0.5 34.0$\pm$3.2 93.5$\pm$2.5 30.0$\pm$2.5 90.5$\pm$3.3 35.1 94.3
FedGIMP 4x 46.3$\pm$3.1 99.6$\pm$0.1 39.5$\pm$3.6 96.4$\pm$1.8 34.7$\pm$2.9 94.7$\pm$2.2 40.2 96.9
8x 41.6$\pm$2.9 98.7$\pm$0.4 34.9$\pm$3.3 94.0$\pm$2.6 30.3$\pm$2.8 90.6$\pm$3.6 35.6 94.4
FedDDA 4x 47.4$\pm$2.8 99.7$\pm$0.1 39.2$\pm$3.4 96.5$\pm$1.5 29.2$\pm$1.4 60.7$\pm$2.1 38.6 85.6
8x 42.9$\pm$3.0 99.1$\pm$0.4 34.1$\pm$3.4 93.9$\pm$2.4 27.8$\pm$2.0 63.5$\pm$2.7 34.9 85.5
FedGAT 4x 50.4$\pm$2.4 99.8$\pm$0.1 39.8$\pm$3.2 96.7$\pm$1.3 36.0$\pm$2.9 95.8$\pm$1.8 42.1 97.4
8x 43.4$\pm$2.9 99.3$\pm$0.4 35.3$\pm$3.1 94.6$\pm$2.1 30.6$\pm$2.8 92.0$\pm$3.1 36.4 95.3
Across-site Site 1 (MoDL5) Site 2 (MoDL3) Site 3 (MoDL7) Average
[…]
Non-FL Single 4x 34.2$\pm$4.5 90.9$\pm$4.4 39.3$\pm$6.9 95.8$\pm$4.1 45.4$\pm$5.0 98.3$\pm$1.8 39.6 95.0
8x 30.8$\pm$4.1 87.8$\pm$5.5 34.5$\pm$6.5 92.8$\pm$5.6 39.0$\pm$4.2 97.1$\pm$2.4 34.8 92.6
Central 4x 39.0$\pm$4.6 96.4$\pm$1.8 40.8$\pm$7.1 96.8$\pm$3.2 47.2$\pm$5.4 98.6$\pm$1.5 42.3 97.3
8x 33.8$\pm$4.5 93.5$\pm$3.2 35.3$\pm$6.2 93.4$\pm$5.3 41.8$\pm$4.4 97.8$\pm$2.0 37.0 94.9
FL methods FedDF 4x 36.4$\pm$3.0 95.2$\pm$1.9 39.3$\pm$2.7 96.2$\pm$1.7 43.4$\pm$3.0 98.1$\pm$1.0 39.7 96.5
8x 31.9$\pm$2.9 92.1$\pm$2.9 34.9$\pm$2.8 93.6$\pm$2.5 37.9$\pm$3.2 96.3$\pm$1.8 34.9 94.0
FedMD 4x 36.3$\pm$2.9 95.2$\pm$1.9 39.6$\pm$2.8 96.3$\pm$1.8 43.4$\pm$3.0 98.1$\pm$1.0 39.8 96.5
8x 31.9$\pm$2.9 92.2$\pm$2.8 34.9$\pm$2.8 93.6$\pm$2.5 37.6$\pm$3.3 96.4$\pm$1.7 34.8 94.1
FedGIMP 4x 37.0$\pm$4.6 94.9$\pm$3.1 38.7$\pm$6.5 95.3$\pm$4.6 44.1$\pm$4.1 98.3$\pm$1.7 39.9 96.2
8x 32.9$\pm$4.5 92.3$\pm$3.9 34.7$\pm$6.2 92.3$\pm$5.8 39.0$\pm$3.7 96.8$\pm$2.8 35.5 93.8
FedDDA 4x 37.5$\pm$4.0 95.5$\pm$2.5 39.8$\pm$6.8 95.9$\pm$4.1 32.3$\pm$1.4 60.0$\pm$6.0 36.5 83.8
8x 33.0$\pm$4.2 92.7$\pm$3.5 35.1$\pm$6.2 93.2$\pm$5.4 31.5$\pm$1.8 59.9$\pm$8.0 33.2 81.9
FedGAT 4x 38.1$\pm$4.5 96.1$\pm$1.9 40.1$\pm$7.0 96.4$\pm$3.6 46.8$\pm$5.3 98.4$\pm$1.9 41.7 97.0
8x 33.4$\pm$4.2 93.2$\pm$3.4 35.3$\pm$6.8 93.8$\pm$5.4 40.4$\pm$4.6 97.4$\pm$2.4 36.4 94.8
Table 4: Within-site and across-site reconstruction performance for multi-coil datasets (Site 1: fastMRI-knee, Site 2: fastMRI-brain, Site 3: UMRAM). Each site employs an MoDL variant with different number of cascades.
Within-site Site 1 (MoDL3) Site 2 (MoDL5) Site 3 (MoDL7) Average
[…]
Non-FL Single 4x 33.8$\pm$2.3 93.4$\pm$1.5 40.3$\pm$3.5 98.0$\pm$0.8 50.0$\pm$2.1 99.6$\pm$0.2 41.4 97.0
8x 29.4$\pm$2.1 83.9$\pm$3.1 34.2$\pm$3.4 94.3$\pm$2.0 42.1$\pm$1.9 98.3$\pm$0.7 35.2 92.2
Central 4x 33.7$\pm$2.3 93.2$\pm$1.6 41.0$\pm$3.8 98.1$\pm$1.2 46.4$\pm$2.2 99.3$\pm$0.3 40.4 96.9
8x 29.4$\pm$2.2 83.6$\pm$3.1 34.3$\pm$3.5 94.1$\pm$2.3 39.7$\pm$1.9 97.9$\pm$0.7 34.5 91.9
FL methods FedDF 4x 32.3$\pm$2.4 90.5$\pm$2.4 38.9$\pm$3.6 95.7$\pm$1.5 44.6$\pm$2.7 98.5$\pm$0.6 38.6 94.9
8x 28.4$\pm$2.3 80.7$\pm$4.1 33.0$\pm$3.7 91.4$\pm$3.4 37.0$\pm$2.3 96.7$\pm$1.4 32.8 89.6
FedMD 4x 32.9$\pm$2.4 92.0$\pm$1.8 39.9$\pm$3.9 97.5$\pm$1.1 44.6$\pm$3.1 99.2$\pm$0.3 39.1 96.3
8x 28.2$\pm$2.4 81.6$\pm$3.5 33.9$\pm$3.8 93.7$\pm$2.6 37.6$\pm$2.6 97.1$\pm$1.2 33.2 90.8
FedGIMP 4x 33.2$\pm$2.3 93.1$\pm$1.5 40.5$\pm$3.4 97.9$\pm$0.9 45.0$\pm$2.1 99.0$\pm$0.3 39.6 96.7
8x 29.2$\pm$2.1 83.6$\pm$3.1 33.5$\pm$3.6 93.8$\pm$2.3 39.5$\pm$1.7 97.7$\pm$0.8 34.1 91.7
FedDDA 4x 32.5$\pm$2.3 91.8$\pm$1.8 39.6$\pm$3.6 97.5$\pm$1.4 42.6$\pm$2.9 97.3$\pm$0.8 38.2 95.5
8x 28.1$\pm$2.2 81.9$\pm$2.9 33.3$\pm$3.5 93.6$\pm$2.8 37.5$\pm$2.2 93.6$\pm$1.5 33.0 89.7
FedGAT 4x 34.1$\pm$2.2 97.3$\pm$1.7 40.9$\pm$3.6 98.0$\pm$1.3 45.1$\pm$2.1 99.2$\pm$0.3 40.0 96.9
8x 29.5$\pm$2.2 83.7$\pm$3.1 34.2$\pm$3.5 93.9$\pm$2.5 40.2$\pm$1.7 97.8$\pm$0.7 34.6 91.8
Across-site Site 1 (MoDL3) Site 2 (MoDL5) Site 3 (MoDL7) Average
[…]
Non-FL Single 4x 38.4$\pm$3.3 97.3$\pm$1.7 39.4$\pm$4.5 96.9$\pm$2.1 37.6$\pm$4.6 96.0$\pm$2.3 38.5 96.7
8x 32.4$\pm$3.0 92.6$\pm$2.9 33.0$\pm$3.6 90.6$\pm$6.1 31.5$\pm$4.0 88.1$\pm$5.4 32.3 90.4
Central 4x 39.7$\pm$3.4 97.6$\pm$1.7 40.6$\pm$4.8 97.4$\pm$2.0 40.4$\pm$4.1 97.5$\pm$1.3 40.2 97.5
8x 34.3$\pm$3.1 93.9$\pm$3.0 33.6$\pm$3.9 91.4$\pm$5.7 34.5$\pm$3.5 91.9$\pm$4.0 34.1 92.4
FL methods FedDF 4x 38.2$\pm$3.4 96.4$\pm$2.1 38.3$\pm$4.6 95.2$\pm$3.6 37.4$\pm$4.1 94.7$\pm$2.4 38.0 95.4
8x 32.9$\pm$3.1 92.4$\pm$3.5 32.7$\pm$4.1 89.4$\pm$7.3 30.2$\pm$3.6 86.8$\pm$5.6 31.9 89.5
FedMD 4x 38.8$\pm$3.4 97.0$\pm$1.7 39.3$\pm$4.8 96.8$\pm$2.3 38.4$\pm$4.5 95.9$\pm$2.3 38.8 96.6
8x 32.5$\pm$2.9 92.2$\pm$3.0 33.2$\pm$3.8 90.9$\pm$6.1 31.9$\pm$3.8 87.5$\pm$5.6 32.5 90.2
FedGIMP 4x 37.6$\pm$3.2 97.2$\pm$1.6 38.9$\pm$4.3 96.8$\pm$2.2 37.4$\pm$4.4 96.4$\pm$1.8 38.0 96.8
8x 32.1$\pm$3.0 92.5$\pm$2.9 32.5$\pm$3.5 90.6$\pm$6.1 31.7$\pm$4.1 89.3$\pm$5.2 32.1 90.8
FedDDA 4x 35.5$\pm$3.3 94.5$\pm$2.4 37.8$\pm$3.8 95.7$\pm$1.5 36.6$\pm$4.2 95.2$\pm$2.4 36.6 95.1
8x 30.2$\pm$3.6 88.7$\pm$4.0 32.1$\pm$3.4 89.2$\pm$5.0 30.8$\pm$3.7 86.3$\pm$5.5 31.0 88.1
FedGAT 4x 39.3$\pm$3.1 97.5$\pm$1.5 39.7$\pm$4.5 97.2$\pm$2.1 39.2$\pm$4.3 97.1$\pm$1.6 39.4 97.3
8x 33.5$\pm$3.0 92.9$\pm$3.1 33.6$\pm$3.9 90.9$\pm$5.8 33.0$\pm$3.9 90.3$\pm$4.6 33.4 91.4